# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v4

@@ 3.5 Geometric Interpretation: KL-Mirror Descent in $u$-Space
Linearize $\mathcal{L}$ at $u_{t}$: $\mathcal{L}(u)\approx\mathcal{L}(u_{t})+\langle g_{t},u-u_{t}\rangle$. A mirror descent step with generator $A$ minimizes $\langle g_{t},u\rangle+\tfrac{1}{\eta}D_{A}(u\|u_{t})$. Using $D_{A}(u\|u_{t})=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$ for the softmax family yields (11). ∎∎
Dual form in $p$-space. Via the Legendre map $p=\nabla A(u)$, the step (11) is equivalent to the entropic proximal in probability space
$\tilde{p}_{t+1}=\arg\min_{p\in\Delta}\ \langle g_{t},p\rangle+\tfrac{1}{\eta}\mathrm{KL}\!\big(p\,\|\,\tilde{p}_{t}\big),$
whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta g_{t})$.
Multiplicative-weights view.
The solution of (11) is equivalent to the multiplicative-weights update in the anchored distribution:
$\tilde{p}_{t+1}\ \propto\ \tilde{p}_{t}\ \odot\ \exp\!\big(-\eta\,g_{t}\big)\ =\ \tilde{p}_{t}\ \odot\ \exp\!\big(-\eta\,(\tilde{p}_{t}-q)\big),$ (12)
followed by renormalization. This exhibits ADPO as a KL-trust-region step in distribution space.
Step-size $\eta$ vs temperature $\tau$ (heuristic).
Anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau$ rescale curvature by $1/\tau^{2}$ (Lemma 3.3). To maintain comparable “trust-region tightness” in the unanchored score space $s$, one may heuristically choose
[…]
This is a practical guideline; a rigorous derivation would require accounting for the chain rule $\partial u/\partial\theta$ when optimizing over shared parameters $\theta$.
@@ 3.6 Dynamic Anchor Stability
Crucially, Lemma 3.3 holds regardless of anchor choice. The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains optimization to a local quadratic region. While the coordinate $u=(s-s^{\mathrm{ref}})/\tau$ depends on the anchor, the local quadratic form at the optimum only depends on $q$; thus the curvature (Fisher metric) is anchor-agnostic, albeit the optimization objective itself changes with the anchor. Common anchor choices include:
  • Fixed anchor: A reference $\pi_{\text{ref}}$ pre-trained and frozen throughout training, providing a stable coordinate origin for consistent target tracking;
  • Dynamic anchor: A policy $\pi_{\theta_{\text{old}}}$ that periodically updates to the current policy (every $K$ epochs via $\pi_{\text{ref}}\leftarrow\pi_{\theta}$), providing an adaptive coordinate system similar to TRPO’s old-policy mechanism;
  • Teacher anchor: A stronger teacher $\pi_{\text{teacher}}$ for offline distillation (typically fixed).
For dynamic anchors, we characterize the monotonic improvement within each fixed-anchor phase:
Theorem 3.6 (Fixed-anchor-phase monotonic descent in $u$-space).
Consider a training phase where the anchor $\pi_{\mathrm{ref}}$ is held fixed. Define $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ and $u=(s-s^{\mathrm{ref}})/\tau$ are the anchored logits. When directly optimizing $\mathcal{L}$ over $u$ (e.g., tabular setting or each group having independent degrees of freedom), $\mathcal{L}$ is strictly convex on the subspace $\mathbf{1}^{\perp}$ (orthogonal to the all-ones vector) with Hessian $\nabla^{2}\mathcal{L}(u)=\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$, which is positive semidefinite on $\mathbf{1}^{\perp}$ with one zero eigenvalue along $\mathbf{1}$.
Strong convexity (local or bounded-domain):
  • Local: In a neighborhood of $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), the Hessian $\nabla^{2}\mathcal{L}(u)=\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$ approaches $\mathrm{Diag}(q)-qq^{\top}$ up to $o(1)$, so there exists $\mu>0$ (depending on $q$) such that $\mathcal{L}$ is $\mu$-strongly convex on $\mathbf{1}^{\perp}$ in this neighborhood.
  • Bounded domain: If the optimization is constrained to a domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ for all $i$ and some $\varepsilon>0$ (i.e., the current anchored distribution $\tilde{p}_{\theta}(u)$ is bounded away from the probability simplex boundary), then $\nabla^{2}\mathcal{L}(u)\succeq\varepsilon\,P_{\mathbf{1}^{\perp}}$ globally in this domain (one may take $c(\varepsilon)=\varepsilon$).
Under either condition, standard gradient descent with appropriate step size $\eta\in(0,1/L]$ (where $L$ is a local smoothness constant) yields linear (geometric) convergence:
$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\;\leq\;(1-\mu\eta)\,(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})),$
implying monotonic decrease in $\mathrm{KL}(q\|\tilde{p}_{\theta})$ within each fixed-anchor phase.
Note: When the anchor is updated (e.g., $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the objective $\tilde{p}_{\theta}$ is re-normalized with the new anchor, so the loss function changes. We do not claim monotonicity across anchor updates; instead, each fixed-anchor phase provides stable descent, and the anchor update adapts the coordinate system for the next phase.
Caveat for parametrized models: In practice, the anchored logits $u$ are derived from shared parameters $\theta$ (e.g., neural networks), and the optimization problem over $\theta$ is typically non-convex. The above results provide geometric intuition for the loss landscape in the anchored distribution space but do not constitute convergence guarantees for parameterized training.
@@
The log-partition function $A(u)$ is strictly convex (its Hessian is the softmax Fisher metric $\mathrm{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$, which is positive semidefinite with null-space along $\mathbf{1}$). The Hessian depends on the current $\tilde{p}_{\theta}(u)$, not directly on $q$. However, near $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$, it coincides with $\mathrm{Diag}(q)-qq^{\top}$ up to higher-order terms, yielding local strong convexity. Alternatively, if $\tilde{p}_{\theta}(u)$ is constrained to $\varepsilon,1-\varepsilon]^{K}$, the minimum eigenvalue is uniformly bounded below. Standard strongly convex optimization theory then guarantees linear convergence. When the anchor changes, $u$ is redefined, so objectives from different phases are not directly comparable. ∎∎
Implication. This theorem explains why dynamic anchors (Section 5.4) maintain stability during online exploration: within each fixed-anchor phase, the implicit Fisher-metric trust region ensures monotonic descent; across anchor updates, the coordinate system adapts to track policy evolution, similar to TRPO’s mechanism. Our experiments (Section 5) reveal task-dependent trade-offs: dynamic anchoring excels at online exploration in noisy environments (adapting coordinates to evolving policies), while fixed anchoring excels at offline distillation from teachers (maintaining a stable coordinate origin throughout). Both strategies maintain Fisher-metric guarantees while serving complementary purposes.
@@
Listwise: $\frac{\partial\mathcal{L}_{\mathrm{group}}^{\mathrm{ref}}}{\partial s_{i}}=\frac{1}{\tau}\big(\tilde{p}_{\theta}(i|S_{x})-q(i|S_{x})\big).$
@@ 5.5 Anchoring to Teacher Policies: Continuous Control Distillation
  • Anchor flexibility: Whether anchoring to old policy (online RL) or teacher policy (offline distillation), the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ automatically constrains updates. This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
@@ 8 Conclusion
Appendix A Fisher Metric Connection and Relation to KD/TRPO
@@
$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})~=~\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta~+~o(\|\delta\|^{2}),\qquad\delta=u-u^{\star}.$ (15)
The matrix $\mathrm{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family (the Hessian $\nabla^{2}A(u^{\star})$), revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space. Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry, with connections to mirror descent and Bregman proximal methods.
@@
Appendix B Proofs and Derivations
B.1 Proof of Proposition 3.1
@@ Appendix B Proofs and Derivations
(iii) Maximum-entropy RL / offline policy projection.
@@
so $s_{a}/\tau\approx Q(s,a)/\alpha+\mathrm{const}$. Choosing $\tau=\alpha$ gives the soft-optimal Boltzmann form $\pi\propto\exp(Q/\alpha)$.
For advantage-weighted updates with $q(a\mid s)\propto\exp(A^{\pi_{\mathrm{old}}}(s,a)/\beta)$ and $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, minimizing $\mathcal{L}_{\mathrm{ADPO}}$ yields
$\tilde{p}_{\theta}(a\mid s)=\mathrm{softmax}_{a}\!\left(\frac{s_{a}-s_{a}^{\mathrm{old}}}{\tau}\right)\propto\exp(A^{\pi_{\mathrm{old}}}(s,a)/\beta),$
so $(s_{a}-s_{a}^{\mathrm{old}})/\tau\approx A^{\pi_{\mathrm{old}}}(s,a)/\beta$, giving $\pi_{\theta}(a\mid s)\propto\pi_{\mathrm{old}}(a\mid s)\exp((\tau/\beta)A^{\pi_{\mathrm{old}}}(s,a))$. Choosing $\tau=\beta$ recovers AWR/AWAC/MPO projections.
@@
$\mathcal{L}(u)-\mathcal{L}(u^{\star})=\tfrac{1}{2}\,\delta^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big]\delta+o(\|\delta\|^{2})=\tfrac{1}{2}\,\mathrm{Var}_{i\sim q}[\delta_{i}]+o(\|\delta\|^{2}).$
@@
$\mathrm{KL}\big(q\|\tilde{p}_{\theta}\big)\approx\frac{1}{2\tau^{2}}\,\mathrm{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\mathbb{E}_{q}[s-s^{\mathrm{ref}}]\big],$
@@
Appendix C Implementation and Training Details
C.1 Noise Generation Procedures
@@
  • Light: SNR = $\frac{\mathrm{Var}[R^{\star}]}{\sigma^{2}}=1.0$, $p_{\text{out}}=0.05$
@@
C.2 Hyperparameters and Training Configuration
@@ C.3 Metrics Definition
Appendix D Additional Experiments and Ablations
D.1 Statistical Significance Tests
@@
D.2 Effect of Reference Initialization
@@
D.3 Temperature Sensitivity (Full Grid)
@@
D.4 KDE-CDF-Logit Transform Details
@@
Appendix E Reproducibility Checklist
@@
  • Noisy observations $\tilde{R}_{i}$ generated according to Section C.

### v5

@@ 3.5 KL-Mirror Descent Interpretation
A mirror descent step with generator $A$ minimizes $\langle g_{t},u\rangle+\tfrac{1}{\eta}D_{A}(u\|u_{t})$. Using the Bregman divergence relation $D_{A}(u\|u_{t})=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$ for the softmax family yields (13). ∎∎
@@ 3.6 Dynamic Anchor Stability
The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ is anchor-agnostic (depends only on $q$), allowing flexible anchor strategies: (i) fixed $\pi_{\text{ref}}$ (frozen, stable origin); (ii) dynamic $\pi_{\theta_{\text{old}}}$ (periodic updates, adaptive like TRPO); (iii) teacher $\pi_{\text{teacher}}$ (offline distillation). We characterize monotonic improvement within each fixed-anchor phase:
Theorem 3.8 (Fixed-anchor-phase monotonic descent).
When the anchor is fixed, $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ is $\mu$-strongly convex on the quotient space and $L$-smooth on any bounded domain where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$. Consequently, gradient descent with step size $\eta\in(0,1/L]$ (or KL-mirror descent with the corresponding step-size condition) enjoys the linear rate
$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\leq(1-\mu\eta)\big(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})\big)$
within that fixed-anchor phase. When the anchor updates ($\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the coordinate system changes—we do not claim monotonicity across anchor updates.
@@
$A(u)$ is strictly convex with Hessian $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p}_{\theta}(u))-\tilde{p}_{\theta}(u)\tilde{p}_{\theta}(u)^{\top}$ (softmax Fisher metric). Near $u^{\star}$ where $\tilde{p}_{\theta}(u^{\star})=q$, this coincides with $\operatorname{Diag}(q)-qq^{\top}$ up to higher-order terms, yielding local strong convexity. In bounded domains where $\tilde{p}_{\theta}(u)_{i}\in[\varepsilon,1-\varepsilon]$ for all $i$, the Fisher metric satisfies $\nabla^{2}\mathcal{L}(u)\succeq\varepsilon\,P_{\mathbf{1}^{\perp}}$ globally in this domain (one may take $c(\varepsilon)=\varepsilon$ in Corollary 3.6). The spectral upper bound of $\nabla^{2}A(u)$ is also bounded in this domain, yielding $L$-smoothness with $L=L(\varepsilon,K)$. ∎∎
Implication. Within each fixed-anchor phase, Fisher-metric trust region ensures monotonic descent; across anchor updates, the coordinate system adapts (like TRPO). Experiments (Section 5) reveal task-dependent trade-offs: dynamic anchors excel at online exploration; fixed anchors excel at offline distillation.
@@
Listwise: $\frac{\partial\mathcal{L}_{\mathrm{group}}^{\mathrm{ref}}}{\partial s_{i}}=\frac{1}{\tau_{\mathrm{anc}}}\big(\tilde{p}_{\theta}(i\mid S_{x})-q(i\mid S_{x})\big).$
@@ 5.5 Anchoring to Teacher Policies: Continuous Control Distillation
  • Anchor flexibility: Whether anchoring to old policy (online RL) or teacher policy (offline distillation), the Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ automatically constrains updates. This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
@@ 8 Conclusion
Appendix A Connection to Offline RL Methods (AWR/AWAC/MPO)
ADPO connects to advantage-weighted reinforcement learning methods at the level of anchored distributions, though not at the policy level. We detail this partial equivalence here.
A.1 Advantage-Weighted Projection in ADPO
When $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (old policy anchor), $q(a|s)\propto\pi_{\text{old}}(a|s)\exp\!\big(A^{\pi_{\text{old}}}(s,a)/\beta\big)$ where $A^{\pi_{\text{old}}}=Q^{\pi_{\text{old}}}-V^{\pi_{\text{old}}}$, and $\tau=\beta$, ADPO’s anchored distribution matches:
$\tilde{p}_{\theta}(a|s)=\mathrm{softmax}\!\left(\frac{\log\pi_{\theta}(a|s)-\log\pi_{\text{old}}(a|s)}{\beta}\right)\;\propto\;\pi_{\text{old}}(a|s)\exp\!\big(A/\beta\big).$
Derivation of the induced policy: By Proposition 3.4, the optimal $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$. Substituting:
$\pi_{\theta}(a|s)$ $\propto\pi_{\text{old}}(a|s)\cdot\big[\pi_{\text{old}}(a|s)\exp(A/\beta)\big]^{\beta}$
$=\pi_{\text{old}}(a|s)\cdot\pi_{\text{old}}(a|s)^{\beta}\cdot\exp(A)$
$=\pi_{\text{old}}(a|s)^{1+\beta}\cdot\exp\!\big(A^{\pi_{\text{old}}}(s,a)\big).$
A.2 Comparison with Classic AWR/AWAC/MPO
Classic AWR/AWAC M-step [24, 25]: Given advantage-weighted samples, the policy update solves
$\max_{\theta}\mathbb{E}_{(s,a)\sim d}\!\big[\exp(A(s,a)/\beta)\log\pi_{\theta}(a|s)\big],$
which (ignoring normalization) yields $\pi_{\theta}(a|s)\propto\exp(A(s,a)/\beta)$ when capacity is sufficient.
MPO E/M-step [26]: The E-step forms $q(a|s)\propto\exp(Q(s,a)/\eta)$ with dual variable $\eta$ (enforcing a KL budget via Lagrangian). The M-step minimizes $\mathrm{KL}(q\|\pi_{\theta})$, yielding $\pi_{\theta}=q$ under sufficient capacity.
A.3 Key Distinction
ADPO’s equivalence is at the $\tilde{p}-$level (anchored distribution):
$\tilde{p}_{\theta}\;\propto\;\pi_{\text{old}}\exp(A/\beta)\qquad\text{(matches AWR target)},$
but the induced policy $\pi_{\theta}$ differs:
$\text{ADPO:}\quad\pi_{\theta}\;\propto\;\pi_{\text{old}}^{1+\beta}\exp(A),\qquad\text{AWR/MPO:}\quad\pi_{\theta}\;\propto\;\pi_{\text{old}}\exp(A/\beta)\text{ or }\pi_{\theta}=q.$
Interpretation: ADPO’s anchored projection induces a multiplicative fusion $\pi_{\mathrm{ref}}\cdot q^{\tau}$, which preserves the geometric structure of anchoring but differs from classic policy-space updates. The power $1+\beta$ arises from the composition: anchoring at $\pi_{\text{old}}$ followed by temperature-scaled projection toward $q$.
When does this matter? In tabular settings or with full expressivity over discrete actions, both formulations can represent the same optimal behavior (via different parametrizations). In function approximation with shared parameters, the difference becomes significant: ADPO’s form provides an implicit trust region via Fisher geometry, while AWR relies on explicit advantage weighting.
Appendix B Fisher Metric Connection and Relation to KD/TRPO
@@
$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})~=~\tfrac{1}{2}\,\delta^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\delta~+~o(\|\delta\|^{2}),\qquad\delta=u-u^{\star}.$ (15)
The matrix $\operatorname{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family (the Hessian $\nabla^{2}A(u^{\star})$), revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space. Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry, with connections to mirror descent and Bregman proximal methods.
@@
Appendix C Proofs and Derivations
C.1 Proof of Proposition 3.2
@@ Appendix C Proofs and Derivations
(iii) Maximum-entropy RL.
@@
so $s_{a}/\tau\approx Q(s,a)/\alpha+\mathrm{const}$. Choosing $\tau=1$ gives $s_{a}\approx Q(s,a)/\alpha$, hence $\pi_{\theta}\propto\exp(Q/\alpha)$, the soft-optimal Boltzmann form.
@@
$\mathcal{L}(u)-\mathcal{L}(u^{\star})=\tfrac{1}{2}\,\delta^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big]\delta+o(\|\delta\|^{2})=\tfrac{1}{2}\,\operatorname{Var}_{i\sim q}[\delta_{i}]+o(\|\delta\|^{2}).$
@@
$\mathrm{KL}\big(q\|\tilde{p}_{\theta}\big)\approx\frac{1}{2\tau^{2}}\,\operatorname{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\mathbb{E}_{q}[s-s^{\mathrm{ref}}]\big],$
@@
Appendix D Implementation and Training Details
D.1 Noise Generation Procedures
@@
  • Light: SNR = $\frac{\operatorname{Var}[R^{\star}]}{\sigma^{2}}=1.0$, $p_{\text{out}}=0.05$
@@
D.2 Hyperparameters and Training Configuration
@@ D.3 Metrics Definition
Appendix E Additional Experiments and Ablations
E.1 Statistical Significance Tests
@@
E.2 Effect of Reference Initialization
@@
E.3 Temperature Sensitivity (Full Grid)
@@
E.4 KDE-CDF-Logit Transform Details
@@
Appendix F Reproducibility Checklist
@@
  • Noisy observations $\tilde{R}_{i}$ generated according to Section D.